OpenStudy (anonymous):

What is the soultion for y= -x - 5 y = 2x + 4 A. 6,7 ( i think it is this one ) B . 6,8 C. -3, -2 D. -6,7 @Phebe

OpenStudy (phebe):

Its not A

OpenStudy (phebe):

because this"y = 2x + 4 " equals to -2

OpenStudy (phebe):

what do you think the answer is?

OpenStudy (anonymous):

C?

OpenStudy (phebe):

Yes

OpenStudy (anonymous):

How many solutions does the equation -3y + 3y +4 = 4 have ? A. NONE B. INFINITELY MANY C.ONE D. FOUR

OpenStudy (phebe):

C

OpenStudy (anonymous):

WHat about 3x + 6 = -1 - 3 + 4x ? Same question.

OpenStudy (anonymous):

@Phebe

OpenStudy (phebe):

im here

OpenStudy (phebe):

x=10